• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 14:22:06 on May 05, 2008

    Entry number 237181

    This entry is a followup to: 237156

    keyword=Controlled access activity (late entry)

    BigBite magnet powered down at start of work.

    LHRS moved to 13.2 degrees.

    BigBite moved to 54 degrees.

    He bottle replaced (almost empty) -- now 2500psi.

    Both He bags filled. Checked flow and bubbler level for flangle.

    Checked Ethane/Argon bottle -- now 300psi.

    BigBite magnet powered back up at end of work.